7. A tablet of Tylenol TM has a mass of 0.611g. It contains 251 mg of its active ingredient, acetaminophen C8H9NO2.

a. What is the mass percent of acetaminophen in the tablet?
b. If the only nitrogen in the tablet exist in as acetaminophen how many grams of nitrogen are present in the tablet?
c. How many atoms of N are in the tablet?

Could you explain how to do it too please

Thank you
Answer #1

Given mass of acetaminophen in the tablet = 251 mg.

Mass of the Tablet = 0.611 g

a. mass percent of acetaminophen =( 251 * 10^-3 / 0. 611) * 100

= 41 percent.

b. In I mole of C8H9NO2 there is 1 mole of nitrogen.

= 251 * 10^-3 / 151 = no of moles of Nitrogen = 1.66 * 10^-3 moles

molar mass of nitrogen = 14

so no of grams of nitrogen = 1.66 * 10^-3 * 14 = 23.27 mg

c. No of moles of hydrogen = 1.66 * 10^-3

No of atoms = 1.66 * 10^-3 * 6.023 * 10^23

= 10^21 atoms of nitrogen.